1. Find the pH of the following solutions. Make sure you express your answer in the correct

number of significant digits. (12 marks)


a) 0.0010 M HClO4 …..__________ b) 2.35 x 10-3 M KOH …___________


c) 4.678 x 10-5 M HNO3..____­_____ d) 12.0 M HCl ...............…__________


e) 0.20 M Ba(OH)2 .....___________ f) 4.30 M H3O+ ...........…...__________


2. Find the [H3O+] in 0.00256 M HBr. (1 mark)


Answer ________________

3. Find the [H3O+] in 0.80 M LiOH. (2 marks)




Answer ________________

4. Find the [H3O+] in a solution with a pH = 3.216 (1 mark)



Answer ________________

5. Find the [H3O+] in 0.45 M BaO (2 marks)



Answer ________________

6. Find the [OH-] in 0.150 M HClO4 (2 marks)





20



Answer ________________



1
Expert's answer
2022-04-12T05:28:04-0400

1) a) HClO4 + H2O = H3O+ + ClO4-

pH = -lg [H3O+] = -lg (0.0010 M) = 3

b) KOH = K+ + OH-

pH = 14 - pOH = 14 - (-lg [OH-]) = 14 - (-lg(2.35 x 10-3 M)) = 11.37

c) HNO3 + H2O = H3O+ + NO3-

pH = -lg [H3O+] = -lg (4.678 x 10-5 M) = 4.33

d) HCl + H2O = H3O+ + Cl-

pH = -lg [H3O+] = -lg (12.0 M) = -1.08

e) Ba(OH)2 = Ba2+ + 2OH-

pH = 14 - pOH = 14 - (-lg [OH-]) = 14 - (-lg(2 x 0.20 M)) = 13.6

f) pH = -lg [H3O+] = -lg (4.30 M) = -0.63

2) HBr + H2O = H3O+ + Br-

[H3O+] = 0.00256 M

3) LiOH = Li+ + OH-

[H3O+] = 10-14 / [OH-] = 10-14 / 0.80 = 1.25 x 10-14 M

4) pH = 3.216

[H3O+] = 10-pH = 10-3.216 = 6.08 x 10-4 M

5) O2- + H2O = 2OH-

[OH-] = 2 x 0.45 = 0.90 M

[H3O+] = 10-14 / [OH-] = 10-14 / 0.90 = 1.11 x 10-14 M

6) HClO4 + H2O = H3O+ + ClO4-

[H3O+] = 0.150 M

[OH-] = 10-14 / [H3O+] = 10-14 / 0.150 = 6.67 x 10-14 M